OpenStudy (anonymous):

point slope form from 4,4 to 6,8

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

slope formula = (y2-y1)/(x2-x1) = (8-4)/(6-4) = 4/2 = 2

OpenStudy (anonymous):

now plug it in to point slope form of y-y1 = m(x-x1) m is the slope